# Beyond the Visible: Multispectral Vision-Language Learning for Earth Observation

### v1

@@ Abstract
Vision-language models for Earth observation (EO) typically rely on the visual spectrum of data as the only model input, thus failing to leverage the rich spectral information available in the multispectral channels recorded by satellites. Therefore, in this paper, we introduce Llama3-MS-CLIP––the first vision-language model pre-trained with contrastive learning on a large-scale multispectral dataset and report on the performance gains due to the extended spectral range. Furthermore, we present the largest-to-date image-caption dataset for multispectral data, consisting of one million Sentinel-2 samples and corresponding textual descriptions generated with Llama3-LLaVA-Next and Overture Maps data. We develop a scalable captioning pipeline, which is validated by domain experts. We evaluate Llama3-MS-CLIP on multispectral zero-shot image classification and retrieval using three datasets of varying complexity. Our results demonstrate that Llama3-MS-CLIP significantly outperforms other RGB-based approaches, improving classification accuracy by +6.77% on average and retrieval performance by +4.63% mAP compared to the second-best model. Our results emphasize the relevance of multispectral vision-language learning. We release the image-caption dataset, code, and model weights under an open-source license.
@@ 1 Introduction
Vision-language models (VLM) have transformed computer vision, enabling powerful zero-shot learning and cross-modal retrieval capabilities [20, 9]. By learning joint representations of images and text, these models generalize across tasks without requiring task-specific training data. However, existing VLMs, such as CLIP [20], are predominantly trained on natural RGB images, which limits their applicability to specialized domains like Earth observation (EO) [29, 26, 15]. Conversely, the effective utilization of multispectral or multichannel input data in VLMs represents an interesting and underexplored research topic in the machine learning community.
@@
EO relies on satellite imagery to monitor environmental changes, urban expansion, and agriculture [31]. To make use of VLMs for such EO applications, researchers proposed a range of domain-specific EO adaptations of CLIP like  [29, 26, 15, 14]. Those models have been trained on up to five million remote sensing images with aligned image captions. Despite demonstrating impressive performance, these models only rely on RGB input channels, instead of leveraging the full spectral range available in multispectral (MS) satellite data for maximum effectiveness. Satellites like Sentinel-2 provide up to 13 spectral bands, capturing rich and important information far beyond visible wavelengths. Yet, up until today, there exists no large-scale multispectral dataset with image captions that is publicly available for the research community and improves the training of VLMs for EO.
To address this gap, we generate approximately one million captions for a popular, multispectral EO dataset derived from the Sentinel satellites called SSL4EO-S12 [25]. For this, we develop an automated captioning approach using metadata tags from Overture maps and the multimodal large language model (MLLM) Llama3-LLaVA-Next-8B [11]. These captions provide semantic grounding for contrastive learning, allowing the model to align multispectral image representations with natural language. We then perform continual pre-training on OpenCLIP [9], adapting the model to the EO domain. Our model, Llama3-MS-CLIP [^1] Built with Meta Llama 3. While the model itself is not based on Llama 3 but OpenCLIP B/16, it is trained on captions generated by a Llama 3-derivative model. Therefore, the model name starts with Llama 3 following its license (https://github.com/meta-llama/llama3/blob/main/LICENSE)., outperforms other VLMs on a range of downstream applications as depicted in Figure 1. The results demonstrate that multispectral data significantly enhances vision-language learning in EO, unlocking capabilities that RGB-based models fail to capture.
Our contributions are threefold: (1) We create the largest multispectral image-caption dataset for EO, (2) we present the first multispectral EO VLM, surpassing current state-of-the-art performance, and (3) we propose novel best practices for model development and image captioning with multispectral data. We release the dataset, code, and model weights under an open-source license.
@@ 2 Related Work
One of the major challenges in EO vision-language learning is the lack of large-scale image-text datasets with multispectral data [27]. Unlike natural images, satellite images do not inherently come with descriptive text. Most EO datasets provide only categorical labels or metadata, making it difficult to train models that require diverse textual supervision. While some approaches have attempted to generate captions for EO images, they often rely on metadata-based descriptions [26] or manually curated annotations [15]. These approaches do not scale to the large volumes of data required to train robust vision-language models. UCMC [18], RSICD [16], RSITMD [28] and Sydney Captions [19] are some famous examples of human-curated EO datasets that are often limited by their small size. RS5M [29] collected five million images from eleven source datasets and used BLIP-2 [12] to generate captions for the RGB images. RSCLIP [14] introduced a pseudo-labeling technique that automatically generates pseudo-labels from unlabeled data. ChatEarthNet [27] is the first larger multi-spectral dataset with over 100k samples, using ChatGPT-3.5 to generate captions. However, the captions are solely based on a small set of land-use/land-cover classes without visual input for the LLM.
Recent approaches have adapted VLMs to EO by continual pre-training on domain-specific RGB datasets. For instance, SkyCLIP [26], RSCLIP [14] and RemoteCLIP [15] built large-scale image-text datasets and adapted CLIP-based backbones. GeoRSCLIP [29] was pre-trained on RS5M, the largest known image-test dataset in the domain with five million images. Differently, GRAFT [17] utilized co-located street-view images to correlate satellite imagery with language. Despite these advancements, all these methods rely solely on RGB data, ignoring the rich spectral information available in multispectral EO imagery.
In contrast to prior work, we introduce a self-supervised approach based on multimodal large language models (MLLM) and Overture annotations, to automatically generate captions for multispectral EO imagery (i.e. Sentinel-2 images). By fine-tuning OpenCLIP on this dataset, we enable multispectral vision-language learning, allowing the model to leverage spectral information beyond the visual spectrum.
@@ 3 Automated Captioning
We generate captions by employing a multimodal large language model, specifically Llama-LLaVA-Next-8B. The model was selected after evaluating three MLLM models, which we outline in the supplementary material in more detail. The process consists of the following steps: First, we extract the RGB channels from S-2 L2A data and scale it to a uint8 value range of 0––255 as no publicly available MLLM supports multispectral inputs. The images are resized to 224 $\times$ 224 pixels as input for the captioning model. We then extract geographical tags from the Overture Maps base layer [^2] Overture Maps: https://docs.overturemaps.org (Version: 2024-03-12-alpha.0) that provides additional contextual information about land cover, infrastructure, and other features in the satellite image. The geographical instances are then sorted and filtered by size, i.e., all features smaller than 2500 square meters (5 $\times$ 5 pixels) are omitted. We use all tags of each instance as they include additional information like intermittent rivers. We further add the names of places to avoid hallucinations. Otherwise, we observe that the model often refers to popular places incorrectly, like labeling most universities as Harvard or Berkeley. Finally, we prompt the MLLM to generate captions in a structured manner by following a chain-of-thought approach: first, the model is prompted to generate three relevant question-answer pairs, which then guide the model in producing the final caption. The prompt includes further instructions to avoid hallucinations and increase the caption quality (see supplementary material). We repeat the generation if any Q&A pair or the caption is missing in the output.
@@
Figure 2 shows example images and their corresponding generated captions. While we do observe several hallucinations in the generated captions, they are also more diverse and include more details than other image-caption datasets such as SkyScript or RS5M which are based on heuristics [26] or the much simpler MLLM BLIP2 [29, 13] (see supplementary material for examples). Figure 2 includes three examples with hallucinations to showcase their different forms. For example, the model sometimes imagines landmasses in ocean patches, man-made or water features, and provides wrong counts or length estimations.
@@
Figure 2: Examples of image-caption pairs of high quality (left) and hallucination examples (right) of our generated pretraining dataset. We highlight hallucinations in red.
@@
We evaluated the captions quantitatively by comparing our validation set with manually labeled EO datasets: RSITMD [28], RSICD [16], and UCM Captions [18]. The generated captions exhibit a much higher average n-gram diversity of 0.75 compared to only 0.48 to 0.49 in the three human-annotated datasets. The similarity between captions is also lower, showing a higher lexical variety in the SSL4EO-S12-captions. We provide the detailed analysis in the supplementary material.
@@
To assess the quality of our dataset, we asked domain experts to conduct a manual evaluation using a random subset with more than one thousand captions from the validation split. Domain experts rated the captions based on completeness and presence of hallucinations. The caption completeness represents if all relevant features in the image are covered in the caption and is measured on a scale from 0 (Terrible) to 5 (Excellent). Additionally, the experts counted hallucinations. Figure 3 presents the distribution of expert ratings, indicating that over 85% of the captions are considered to include most of the relevant features in the image. The human assessment further demonstrates that two thirds of the evaluated data is free from hallucinations. If hallucinations are present, we typically observe only one hallucinated feature within an image.
The manual evaluation of the generated captions demonstrates that automated captioning with a general-purpose MLLM is feasible and leads to a diverse set of captions. Different from datasets like ChatEarthNet [27] or SkyScript [26] that do not use multimodal LMMs, our pipeline can capture scene-specific features like snow, clouds, or colors. While we do want to highlight the challenge of hallucinations in the dataset, our experiments show that VLMs can learn semantic concepts from the correct annotations. Furthermore, the alignment with S-1 GRD data in SSL4EO-S12 and the question-answer pairs provide additional potential for the EO community.
@@ 4 Llama3-MS-CLIP
Llama3-MS-CLIP is trained with self-supervised contrastive language-image pre-training (CLIP) [20], visualized in Figure 4. We modified the input layer to handle Sentinel-2’s spectral bands beyond RGB by extending the patch embedding for the additional channels. We initialize the corresponding weights with zero tensors so that, during continual pretraining, the model starts from RGB input and can iteratively include additional channels based on optimizing the loss landscape. Hence, the model can slowly learn to leverage the additional information. Our initial experiments suggest that this initialization strategy outperforms random initialization, where the continual pretraining would be disrupted due to the noise that originates from the random weights for the additional channels.
@@
Following CLIP, we utilize the InfoNCE loss, a contrastive loss function, to align embeddings of semantically similar samples, while separating those that are semantically dissimilar through cross-modal supervision. The InfoNCE loss encourages the embeddings of matching (positive) pairs $(x_{i},y_{i})$ to be similar, while pushing apart non-matching (negative) pairs $(x_{i},y_{j})$ with $j\neq i$. Here, $\mathrm{sim}(\cdot,\cdot)$ is a pairwise similarity measure, and $\tau$ is a temperature parameter that scales the logits. Minimizing the loss thus maximizes the similarity of each positive pair relative to all negative pairs. We summarize the training objective in equation 1.
@@ 5 Experimental Setup
Pre-training. We use the implementation and model weights provided by OpenCLIP [9] and performed continual pre-training of the ViT-B/16 using the SSL4EO-S12-captions. The model has 150 million parameters, split between the image and text encoder and was initially pre-trained on LAION-2B, an English subset of the LAION-5B [22] dataset. While many EO models use the ViT-B/32 version with a patch size of 32 [29, 26, 15], we find a patch size of 16 more appropriate for our low-resolution images with many details. We used an AdamW optimizer with a learning rate of 4e-5, 50 warm-up steps and a cosine decay scheduler that updated after each training step. The model was trained for up to five epochs on NVIDIA-A100 GPUs with a global batch size of 1200. The final model was selected based on the lowest validation loss, which was reached after three epochs.
@@
Benchmark Models. We evaluate OpenCLIP [5] ViT-B/16 and ViT-L/14 as well as three RGB-based EO-specific models based on ViT-B/32 backbones. SkyCLIP [26] used remote sensing images with rich semantics covered in Open Street Map to construct a dataset comprising 2.6 million images and generated captions with a simple heuristic by just listing all Open Street Map tags. They performed fully unfrozen continual pre-training using the ViT B/32 backbone initialized from the LAION 2B weights by OpenCLIP [9]. RemoteCLIP [15] proposed a data scaling approach to existing datasets via annotation unification. For images with bounding box annotations, a box-to-caption generation approach was applied. For datasets with available semantic segmentation, the mask-to-box conversion method was used to generate captions. The resulting high resolution dataset consisted of 165k images with each image accompanied by five captions. RemoteCLIP is based on the OpenAI CLIP weights [20] and was adapted with fully unfrozen weights. The authors of GeoRSCLIP [29] used the images from BigEarthNet [23] and ten other datasets. They generated captions based on the annotations and metadata using BLIP-2 [12]. Subsequently, they fine-tuned the OpenAI ViT B/32 and ViT H-14 model applying parameter-efficient fine-tuning techniques.
@@
Downstream Datasets. Our zero-shot evaluation focuses on low-resolution Sentinel-2 imagery from EuroSAT [8], BigEarthNet [23], and METER-ML [30]. EuroSAT includes 64 $\times$ 64 patches from ten land-use/land-cover (LULC) classes. BigEarthNet consists of S-2 L2A patches with 19 multi-labels covering a more diverse set of LULC classes and has an input size of 120 $\times$ 120 pixels. Finally, METER-ML covers images with seven classes of different methane sources like landfills, coal mines, or natural gas processing plants. METER-ML includes S-2 images of size 72 $\times$ 72 as well as high resolution RGB images from NAIP with size 720 $\times$ 720. Methane is visible in the SWIR S-2 bands (band 11 and 12) and therefore an especially interesting downstream task. Because EuroSAT and METER-ML only include S-2 L1C data, we downloaded L2A data for their test sets to better align the inputs with the pre-training data. We observed improvements for METER-ML and therefore evaluated on L2A data for this task. We perform additional experiments with the METER-ML-NAIP data as well as with the RESISC45 [4] dataset. The latter includes RGB images of size 256 $\times$ 256 with a spatial resolution ranging from 30 m to 0.2 m. The 45 scene classes range from landscapes like wetland to large objects like airplane.
@@
For the multi-label dataset BigEarthNet, we transform each class into a binary classification task. For each class, we calculate the similarity between the image embedding and the respective text embedding of that class and compare it to the mean similarity between the image and the all other classes, as formulated in Equation 2. Here, $\hat{y}_{i}$ is the predicted label for class $i$, $x$ is the image embedding, $c_{i}$ is the class embedding, $\mathrm{sim}(\cdot,\cdot)$ is the dot-product similarity, and $K$ is the total number of classes. We also compare this method to a negative-class approach (e.g., “other features”), which boosts accuracy but substantially lowers recall and the F1 score (results in the supplementary material).

### v2

@@ Abstract
Vision-language models for Earth observation (EO) typically rely on the visual spectrum of data as the only model input, thus failing to leverage the rich spectral information available in the multispectral channels recorded by satellites. Therefore, we introduce Llama3-MS-CLIP––the first vision-language model pre-trained with contrastive learning on a large-scale multispectral dataset and report on the performance gains due to the extended spectral range. Furthermore, we present the largest-to-date image-caption dataset for multispectral data, consisting of one million Sentinel-2 samples and corresponding textual descriptions generated using Llama3-LLaVA-Next and Overture Maps data. We develop a scalable captioning pipeline, which is validated by domain experts. We evaluate Llama3-MS-CLIP on multispectral zero-shot image classification and retrieval using three datasets of varying complexity. Our results demonstrate that Llama3-MS-CLIP significantly outperforms other RGB-based approaches, improving classification accuracy by +6.77% on average and retrieval performance by +4.63% mAP compared to the second-best model. Our results emphasize the relevance of multispectral vision-language learning. The image-caption dataset, code, and model weights are available at https://github.com/IBM/MS-CLIP.
@@ 1 Introduction
Vision-language models (VLM) have transformed computer vision, enabling powerful zero-shot learning and cross-modal retrieval capabilities [20, 10]. By learning joint representations of images and text, these models generalize across tasks without requiring task-specific training data.However, existing VLMs, such as CLIP [20], are predominantly trained on natural RGB images, limiting their applicability to specialized domains such as Earth observation(EO) [28, 25, 16]. Conversely, effectively utilizing multispectral input data in VLMs represents an interesting and underexplored research topic in the machine learning community.
@@
EO relies on satellite imagery to monitor environmental changes, urban expansion, and agriculture [30]. To use VLMs for such EO applications, researchers proposed a range of domain-specific EO adaptations of CLIP [28, 25, 16, 15]. Those models have been trained on up to five million remote sensing images with aligned image captions. Despite impressive performance, these models rely only on RGB input channels instead of leveraging the full spectral range available in multispectral (MS) satellite data for maximum effectiveness. Satellites like Sentinel-2 provide up to 13 spectral bands, capturing rich information far beyond visible wavelengths. Yet, until today, no large-scale multispectral dataset with image captions is publicly available for the research community.
To address this gap, we generate approximately one million captions for a popular, multispectral EO dataset derived from the Sentinel satellites called SSL4EO-S12 [24]. We develop an automated captioning approach using metadata tags from Overture maps and the multimodal large language model (MLLM) Llama3-LLaVA-Next-8B [12]. These captions provide semantic grounding for contrastive learning, allowing the model to align multispectral image representations with natural language. We then perform continual pre-training on OpenCLIP [10], adapting the model to the EO domain. Our model, Llama3-MS-CLIP [^1] Built with Meta Llama 3. While the model itself is not based on Llama 3 but OpenCLIP B/16, it is trained on captions generated by a Llama 3-derivative model. Therefore, the model name starts with Llama 3 following its license (https://github.com/meta-llama/llama3/blob/main/LICENSE)., outperforms other VLMs on a range of downstream applications as depicted in Figure 1. The results demonstrate that multispectral data significantly enhances vision-language learning in EO, unlocking capabilities that RGB-based models fail to capture.
Our contributions are threefold: (1) We create the largest multispectral image-caption dataset for EO, (2) we present the first multispectral EO VLM, surpassing current state-of-the-art performance, and (3) we propose novel best practices for model development and image captioning with multispectral data. The dataset, code, and model weights are available at https://github.com/IBM/MS-CLIP under a permissive license.
@@ 2 Related Work
One of the major challenges in EO vision-language learning is the lack of large-scale image-text datasets with multispectral data [26]. Unlike natural images, satellite images do not inherently come with descriptive text. Most EO datasets provide only categorical labels or metadata, making it challenging to train models that require diverse textual supervision. While some approaches have attempted to generate captions for EO images, they often rely on metadata-based descriptions [25] or manually curated annotations [16]. These approaches do not scale to the large volumes of data required to train robust vision-language models. UCMC [19], RSICD [17], and RSITMD [27] are some famous examples of human-curated EO datasets that are often limited by their small size. RS5M [28] collected five million images from eleven source datasets and used BLIP-2 [13] to generate captions for the RGB images. RSCLIP [15] introduced a pseudo-labeling technique that automatically generates pseudo-labels from unlabeled data. ChatEarthNet [26] is the first multispectral dataset with over 100k samples, using ChatGPT-3.5 to generate captions. However, the captions are solely based on a small set of land-cover classes without visual input for the LLM.
Recent approaches have adapted VLMs to EO by continual pre-training on domain-specific RGB datasets. For instance, SkyCLIP [25], RSCLIP [15], and RemoteCLIP [16] built large-scale image-text datasets and adapted CLIP-based backbones. GeoRSCLIP [28] was pre-trained on RS5M, the largest known image-test dataset in the domain with five million images. GRAFT [18] utilized co-located street-view images to correlate satellite imagery with language. Despite these advancements, all these methods rely solely on RGB data, ignoring the rich spectral information available in multispectral EO imagery.We further note the emergence of autoregressive approaches that generate language output based on optical or radar images, like in TerraMind [4].
In contrast to prior work, we introduce a self-supervised approach based on multimodal large language models (MLLM) and Overture annotations to automatically generate captions for multispectral EO imagery (i.e., Sentinel-2 images). By fine-tuning OpenCLIP on this dataset, we enable multispectral vision-language learning, allowing the model to leverage spectral information beyond the visible spectrum.
@@ 3 Automated Captioning
We generate captions by employing a multimodal large language model, specifically Llama-LLaVA-Next-8B. The model was selected based on a qualitative comparison and a quantitative evaluation of three MLLM models using METEOR (Metric for Evaluation of Translation with Explicit Ordering) [8], comparing ground truth captions with generations. We tested BLIP2 [13], used for the captioning in the RS5M [28] dataset, Llama3-LLaVA-Next-8B [12], and RS-LLaVA [1], a domain-specific adaptation of LLaVA 1.5. We assessed these models using UCM Captions [19], RSICD [17], and RSITMD [27] that provide human-annotated captions. Llama3-LLaVA-Next-8B reaches an average METEOR score of 0.20 compared to 0.16 for RS-LLaVA and 0.10 for BLIP2. All scores and some examples are provided in the supplementary material.
@@
The captioning process consists of the following steps: First, we extract the RGB channels from S-2 L2A data and scale it to a uint8 value range of 0––255 as no publicly available MLLM supports multispectral inputs. The images are resized to 224 $\times$ 224 pixels as input for the captioning model. We then extract geographical tags from the Overture Maps base layer [^2] Overture Maps: https://docs.overturemaps.org (Version: 2024-03-12-alpha.0) that provides additional contextual information about land cover, infrastructure, and other features in the satellite image. The geographical instances are then sorted and filtered by size, i.e., all features smaller than 2500 square meters (5 $\times$ 5 pixels) are omitted. We use all tags of each instance as they include additional information, like intermittent rivers. We further add the names of places to avoid hallucinations. Otherwise, we observe that the model often refers to popular places incorrectly, like labeling most universities as Harvard or Berkeley. Finally, we prompt the MLLM to generate captions in a structured manner by following a chain-of-thought approach. First, the model is prompted to generate three relevant question-answer pairs, guiding the model in producing the final caption. The prompt includes further instructions to avoid hallucinations and increase the caption quality. We repeat the generation if any Q&A pair or the caption is missing in the output. We include the prompt and other details on the captioning process in the supplementary material.
@@
Figure 2 shows example images and their corresponding generated captions. While we do observe several hallucinations in the generated captions, they are also more diverse and include more details than other image-caption datasets such as SkyScript or RS5M, which are based on heuristics [25] or the much simpler MLLM BLIP2 [28, 14] (see supplementary material for examples). Figure 2 includes three examples with hallucinations to showcase their different forms. For example, the model sometimes imagines landmasses in ocean patches, man-made or water features, and provides wrong counts or length estimations.
@@
Figure 2: Examples of image-caption pairs of high quality (left) and hallucination examples (right) of our generated pre-training dataset. We highlight hallucinations in red.
@@
We evaluate the captions quantitatively by comparing our validation set with manually labeled EO datasets: RSITMD [27], RSICD [17], and UCM Captions [19]. The generated captions exhibit a much higher average n-gram diversity of 0.75 compared to only 0.48 to 0.49 in the three human-annotated datasets. The similarity between captions is also lower, showing a higher lexical variety in the SSL4EO-S12-captions.
@@
To assess the quality of our dataset, we asked domain experts33314 researchers from the FAST-EO project working in the Earth observation domain. to conduct a manual evaluation using a random subset with more than one thousand captions from the validation split. Domain experts rated the captions based on completeness and presence of hallucinations. The caption completeness represents whether all relevant features in the image are covered in the caption and is measured on a scale from 0 (Terrible) to 5 (Excellent). Additionally, the experts counted hallucinations. Figure 3 presents the distribution of expert ratings, indicating that over 85% of the captions are considered to include most of the relevant features in the image. The human assessment further demonstrates that two-thirds of the evaluated data is free from hallucinations. If hallucinations are present, we typically observe only one hallucinated feature within an image. We provide details of the quantitative comparison and the human evaluation in the supplementary material.
The manual evaluation of the generated captions demonstrates that automated captioning with a general-purpose MLLM and additionally provided tags is feasible and leads to mostly correct captions. Furthermore, the quantitative assessment of the full validation set indicates that the captioning model uses a more diverse vocabulary than existing human-annotated datasets. Different from datasets like ChatEarthNet [26] or SkyScript [25] that do not use multimodal LMMs, our pipeline can capture scene-specific features like snow, clouds, or colors. While we do want to highlight the challenge of hallucinations in the dataset, our experiments show that VLMs can learn semantic concepts from the correct annotations. Furthermore, the alignment with S-1 GRD data in SSL4EO-S12 and the question-answer pairs provides additional potential for the EO community.
@@ 4 Llama3-MS-CLIP
Llama3-MS-CLIP is trained with self-supervised contrastive language-image pre-training (CLIP) [20], visualized in Figure 4. We modified the input layer to handle Sentinel-2’s spectral bands beyond RGB by extending the patch embedding for the additional channels. We initialize the corresponding weights with zero tensors so that during continual pre-training, the model starts from RGB input and can iteratively include additional channels based on optimizing the loss landscape. Hence, the model can slowly learn to leverage the additional information. Our initial experiments suggest that this initialization strategy outperforms random initialization, where the continual pre-training would be disrupted due to the noise that originates from the random weights for the additional channels.
@@
Following CLIP, we utilize the InfoNCE loss, a contrastive loss function, to align embeddings of semantically similar samples while separating semantically dissimilar ones through cross-modal supervision. The InfoNCE loss encourages the embeddings of matching (positive) pairs $(x_{i},y_{i})$ to be similar, while pushing apart non-matching (negative) pairs $(x_{i},y_{j})$ with $j\neq i$. Here, $\mathrm{sim}(\cdot,\cdot)$ is a pairwise similarity measure, and $\tau$ is a temperature parameter that scales the logits. Minimizing the loss thus maximizes the similarity of each positive pair relative to all negative pairs. We summarize the training objective in equation 1.
@@ 5 Experimental Setup
Pre-training. We use the implementation and model weights provided by OpenCLIP [10] and perform continual pre-training of the ViT-B/16 using the SSL4EO-S12-captions. The model has 150 million parameters, split between the image and text encoders, and was initially pre-trained on LAION-2B, an English subset of the LAION-5B [21] dataset. While many EO models use the ViT-B/32 version with a patch size of 32 [28, 25, 16], we find a patch size of 16 more appropriate for low-resolution images with many details. We used an AdamW optimizer with a learning rate of 4e-5, 50 warm-up steps, and a cosine decay scheduler that updated after each training step. The model was trained for up to five epochs on NVIDIA-A100 GPUs with a global batch size of 1200. The final model was selected based on the lowest validation loss reached after one epoch for the multispectral version and two epochs for the RGB version. Based on prior experiments, all layers are unfrozen during the pre-training of Llama3-MS-CLIP, but only the projection layers are trained for the RGB data.
@@
Benchmark Models. We evaluate OpenCLIP [6] ViT-B/16 and ViT-L/14 as well as three RGB-based EO-specific models based on ViT-B/32 backbones. SkyCLIP [25] used remote sensing images with rich semantics covered in Open Street Map to construct a dataset comprising 2.6 million images and generated captions with a simple heuristic by just listing all Open Street Map tags. They performed fully unfrozen continual pre-training using the ViT B/32 backbone initialized from the LAION 2B weights by OpenCLIP [10]. RemoteCLIP [16] proposed a data scaling approach to existing datasets via annotation unification. For images with bounding box annotations, a box-to-caption generation approach was applied. The mask-to-box conversion method was used to generate captions for datasets with available semantic segmentation. The resulting high-resolution dataset consisted of 165k images, each accompanied by five captions. RemoteCLIP is based on the OpenAI CLIP weights [20] and was adapted with fully unfrozen weights. The authors of GeoRSCLIP [28] used the images from BigEarthNet [22] and ten other datasets. They generated captions based on the annotations and metadata using BLIP-2 [13]. Subsequently, they fine-tuned the OpenAI ViT B/32 and ViT H/14 models applying parameter-efficient fine-tuning techniques.
@@
Downstream Datasets. Our zero-shot evaluation focuses on low-resolution Sentinel-2 imagery from EuroSAT [9], BigEarthNet [22], and METER-ML [29]. EuroSAT includes 64 $\times$ 64 patches from ten land-use/land-cover (LULC) classes. BigEarthNet consists of S-2 L2A patches with 19 multi-labels covering a more diverse set of LULC classes and has an input size of 120 $\times$ 120 pixels. Finally, METER-ML covers images with seven classes of different methane sources like landfills, coal mines, or natural gas processing plants. METER-ML includes S-2 images of size 72 $\times$ 72 and high-resolution RGB images from NAIP with size 720 $\times$ 720. Methane is visible in the SWIR S-2 bands (bands 11 and 12) and, therefore, is an especially interesting downstream task. Since EuroSAT and METER-ML only include S-2 L1C data, we downloaded L2A data for their test sets to better align the inputs with the pre-training data. We observed improvements for METER-ML and therefore evaluated on L2A data for this task. We perform additional experiments with the METER-ML-NAIP data and the RESISC45 [5] dataset. The latter includes RGB images of size 256 $\times$ 256 with a spatial resolution ranging from 30 m to 0.2 m. The 45 scene classes range from landscapes like wetland to large objects like airplane.
@@
For the multi-label dataset BigEarthNet, we transform each class into a binary classification task. For each class, we calculate the similarity between the image embedding and the respective text embedding of that class and compare it to the mean similarity between the image and all other classes, as formulated in Equation 2. Here, $\hat{y}_{i}$ is the predicted label for class $i$, $x$ is the image embedding, $c_{i}$ is the class embedding, $\mathrm{sim}(\cdot,\cdot)$ is the dot-product similarity, and $K$ is the total number of classes. We also compare this method to a negative-class approach (e.g., "other features"), which boosts accuracy but substantially lowers recall and the F1 score (results in the supplementary material).